What are the equations, in slope-intercept form, of the lines that have a slope of 3/5 and form a triangle of 30 square units with the x- and y- axis?
Found 2 solutions by santity, MathTherapy:
Answer by santity(1)


using y=mx+c,where y=30,x=3o,m=3/5 ,
then,30.30=3/5.30.30+c,
Where c is the slope intercept
900=3.900/5+c
900=2,700/5+c
900=540+c
C=900-540=360

Answer by MathTherapy(10552)


What are the equations, in slope-intercept form, of the lines that have a slope of 3/5 and form a triangle of 30 square units with the x- and y- axis?